# Augmenting Molecular Graphs with Geometries via Machine Learning Interatomic Potentials

### v1

@@ 3 Related Work
Molecular Representation Learning. Learning informative molecular representations is critical for accurately predicting molecular properties. Early efforts focused on pre-training models on SMILES strings using language modeling techniques such as BERT (Devlin et al., 2019), exemplified by SMILES-BERT (Wang et al., 2019). Subsequently, attention shifted toward pre-training on 2D molecular graphs by designing self-supervised learning tasks (Hu et al., 2019; Rong et al., 2020). Other works further explored learning shared representations between 2D and 3D molecular graphs through contrastive learning (Liu et al., 2021; Stärk et al., 2022). More recently, denoising pre-training (Zaidi et al., 2022; Feng et al., 2023; Ni et al., 2023; Liao et al., 2024) has emerged as a highly effective approach for molecular representation learning. In this paradigm, 3D molecular structures are perturbed with specific noise, and the model is trained to predict the applied noise. Denoising pre-training has been shown to outperform previous pre-training methods, as it is mathematically equivalent to learning an underlying interatomic potential, leading to richer and more physically grounded representations.
Molecular Foundation Models. Inspired by the remarkable success of foundation models in computer vision (Bommasani et al., 2021; Liu et al., 2024; Zhai et al., 2022; Dehghani et al., 2023) and natural language processing (Brown et al., 2020; Touvron et al., 2023; Achiam et al., 2023; Bai et al., 2023), molecular foundation models have also begun to attract significant attention. However, due to the lack of large-scale 3D molecule data with energy and force labels, existing molecule foundation models primarily focus on 1D SMILES strings or 2D molecular graph representations. For example, ChemFM (Cai et al., 2024) adopts self-supervised causal language modeling on SMILES, while MolE (Méndez-Lucio et al., 2022) adapts DeBERTa (He et al., 2020) for molecular graphs. MolFM (Luo et al., 2023) jointly learns molecular representations from graphs, biomedical texts, and knowledge graphs. Nevertheless, none of these models address the fundamental task of learning 3D molecular energies and forces, which limits their applicability to geometry-dependent tasks such as 3D molecular property prediction, geometry optimization, etc.
@@ 4 Experiments
In this section, we demonstrate the effectiveness of the MLIP foundation model across several downstream tasks. In Section 4.1, we benchmark several backbone model candidates on PubChemQCR-S. In Section 4.2, we present geometry optimization using the foundation model. In Section 4.3, we show that foundation model-relaxed geometries can improve molecular property prediction when stable 3D structures are unavailable in the test set. In Section 4.4, we fine-tune the foundation model for 3D molecular property prediction.
@@
Model Energy MAE (meV/atom) Force RMSE (meV/Å)
SchNet $5.30$ $56.55$
PaiNN $5.13$ $46.34$
NequIP $7.37$ $54.78$
SevenNet $8.77$ $47.63$
Allegro $10.86$ $60.71$
FAENet $7.28$ $60.24$
MACE $7.54$ $51.46$
PACE $6.24$ $50.54$
Equiformer $\mathbf{4.69}$ $\mathbf{34.67}$
@@ 4.1 Backbone Models
To select the backbone architecture for the MLIP foundation model, we benchmark several representative MLIP models on our curated dataset. The methods include SchNet (Schütt et al., 2018), FAENet (Duval et al., 2023), NequIP (Batzner et al., 2022), Equiformer (Liao and Smidt, 2022), SevenNet (Park et al., 2024), Allegro (Musaelian et al., 2023), PaiNN (Schütt et al., 2021), PACE (Xu et al., 2024), and MACE (Batatia et al., 2022).
For training efficiency, we curated a smaller subset for model benchmarking, named PubChemQCR-S, which contains 40,979 trajectories and 1,504,431 molecular snapshots from the DFT stage calculations. The benchmark results are summarized in Table 2. In selecting a foundation model backbone, we consider both predictive performance and computational efficiency. Among the candidates, PaiNN demonstrates relatively strong energy prediction accuracy and the second-best force prediction performance on the benchmark while maintaining computational efficiency, making it a strong choice for large-scale pre-training.
@@ 4.2 Geometry Optimization
Table 3: Geometry optimization performance of the MLIP foundation model pre-trained on the curated PubChemQCR dataset.
@@
Metrics. Since the foundation model is trained using data from the DFT relaxation stage, we evaluate its geometry optimization performance by relaxing molecules starting from the first snapshot of the DFT trajectory. The evaluation metrics we adopt are partially based on those proposed in (Tsypin et al., 2023), and include: (1) Average Energy Minimization Percentage, $\overline{\mathrm{pct}}_{T}$, quantifies how much energy is minimized by the MLIP-based optimization relative to the DFT-based optimization; (2) Chemical Accuracy Success Rate, $\mathrm{pct}_{\text{success}}$, measures the percentage of relaxed molecules whose residual energy is within chemical accuracy (commonly defined as 1 kcal/mol); (3) Divergence Rate, $\mathrm{pct}_{\text{div}}$, represents the percentage of relaxed molecules for which either the single-point DFT energy calculation failed or the relaxed DFT energy is higher than the initial energy; (4) Force Convergence Rate, $\mathrm{pct}_{\text{FwT}}$, measures the percentage of relaxed molecules whose maximum force is below a threshold of 0.05 eV/Å. More details can be found in Section A.2.
Results. The results are presented in Table 3. While the value of $\overline{\mathrm{pct}}_{T}$ indicates that the MLIP-relaxed geometries reduce a certain amount of energy compared to the initial structures, a notable portion of conformers remains outside the optimal region. Consequently, the values of $\mathrm{pct}_{\text{success}}$ and $\mathrm{pct}_{\text{FwT}}$ are relatively low. This highlights the inherent challenge of optimizing molecular geometries that are already near the energy minimum—a regime where achieving further relaxation requires extremely accurate modeling of the potential energy surface.
@@
4.3 Molecular Property Prediction with MLIP Foundation Model Relaxed Geometries
Task. As discussed in Section 1, obtaining accurate 3D geometries is crucial for achieving strong performance in quantum property prediction. However, acquiring ground-state geometries typically requires time-consuming DFT-based relaxation. To address this limitation, it is highly desirable to leverage machine learning interatomic potentials (MLIPs) to generate approximate stable geometries that can support downstream property prediction. In this task, we focus on a practical setting where ground-truth 3D geometries are available during training, but only non-stable geometries are available at test time—consistent with the setup used in Uni-Mol+.
Dataset. We use the Molecule3D dataset and predict the HOMO-LUMO gap, a key quantum property of molecular electronic structure. We use a subset of the dataset containing 600,000 molecules for random and scaffold splits. More details about the dataset can be found in Section A.3.
Results. We compare geometry fine-tuning using input structures with varying levels of geometric accuracy. Specifically, we evaluate downstream performance using: (1) RDKit-generated geometries, (2) geometries after the PM3 and HF optimization, i.e. geometries before the DFT relaxation, and (3) MLIP foundation model-relaxed geometries. We include (2) because the foundation model is trained on DFT substage trajectories and the relaxation begins from the first snapshot of the DFT substage. For a fair comparison with Uni-Mol+, we also evaluate Uni-Mol+ using both RDKit-generated geometries and geometries after the PM3 and HF optimization. The results of geometry fine-tuning on the random and scaffold splits are shown in Table 4. These results demonstrate that using foundation model-relaxed geometries consistently improves downstream property prediction. Furthermore, compared to Uni-Mol+, our approach—combining MLIP-based relaxation with a downstream 3D GNN—achieves superior performance, suggesting that this modular pipeline is more effective than architectures specifically designed to bridge the gap between non-stable and ground truth molecular structures.
@@ 4.3 Molecular Property Prediction with MLIP Foundation Model Relaxed Geometries
4.4 Fine-Tuning MLIP Foundation Model for Molecular Property Prediction
Table 5: Results of HOMO-LUMO gap prediction on the $\nabla^{2}$ dataset. Best results are shown in bold, and second-best results are underlined.
@@ 4.4 Fine-Tuning MLIP Foundation Model for Molecular Property Prediction
Task. To demonstrate the effectiveness of the MLIP foundation model for 3D molecular property prediction, we fine-tune the pre-trained foundation model on downstream tasks where the goal is to predict molecular properties given 3D molecular structures. In this setting, ground-truth geometries are used as input. We evaluate the performance on two benchmark datasets, as described below.
Datasets. We use Molecule3D subset created in Section 4.3 and $\nabla^{2}$ (Khrabrov et al., 2024) datasets. The details of the datasets can be found in Section A.3. In this task, we predict the HOMO-LUMO gap, and ground truth 3D geometries are provided. Additional results on the full Molecule3D dataset are provided in Section A.5.1.
Baselines. We select a set of representative models as baselines, including GIN-virtual (Hu et al., 2021), SchNet (Schütt et al., 2017), SphereNet (Liu et al., 2022), ComENet (Wang et al., 2022), DimeNet++ (Gasteiger et al., 2020), TorchMD-Net (Thölke and De Fabritiis, 2022), Uni-Mol+ (Lu et al., 2023), and PaiNN (Schütt et al., 2021). For the Molecule3D dataset, we include 2D models, 3D models, and hybrid approaches like Uni-Mol+. For the $\nabla^{2}$ dataset, we focus exclusively on 3D models.
@@
Results. The results for the Molecule3D random split and scaffold split are reported in Table 6. The fine-tuned foundation model achieves the best performance in both settings. The scaffold split is a more challenging evaluation setup, as it requires models to generalize to out-of-distribution molecular scaffolds. As expected, all methods perform worse on the scaffold split compared to the random split, and the gap between validation and test accuracy is also larger in the scaffold setting. GIN-virtual performs the worst among all methods, as it only uses the 2D molecular graph as input, while the HOMO-LUMO gap is highly sensitive to 3D molecular geometry. Uni-Mol+ outperforms GIN-virtual by incorporating ground-truth 3D geometries during training and learning to predict stable 3D structures from RDKit-initialized conformers. However, it still lags behind 3D GNN models, which use accurate 3D geometries during both training and inference. The results on the $\nabla^{2}$ dataset are presented in Table 5. Again, the fine-tuned foundation model achieves the best performance among the 3D baselines, further demonstrating its effectiveness in downstream molecular quantum property prediction tasks.
Table 6: Results of HOMO-LUMO gap prediction on two splits of the Molecule3D dataset. Force2Prop w/ PaiNN denotes the MLIP foundation model pre-trained on our PubChemQCR dataset. Best results are shown in bold, and second-best results are underlined.
@@
Analysis. Inspired by Zaidi et al. (2022), we conduct an experiment to evaluate the usefulness of pre-trained features. Specifically, we freeze the backbone and fine-tune only the prediction head, and compare it to a baseline where the backbone is randomly initialized and only the prediction head is trained. The training curves are shown in Fig. 4. As expected, fine-tuning only the prediction head performs worse than fine-tuning the entire model, but still significantly outperforms training the head on top of random features. This indicates that pre-trained features provide a meaningful representation for the downstream task. However, to fully adapt the model and achieve optimal performance, it is necessary to fine-tune the entire network.
Figure 5: Results of fine-tuning the MLIP foundation model on Molecule3D with different sizes of downstream data for LABEL:sub@fig:different_downstream_data_size_random random and LABEL:sub@fig:different_downstream_data_size_scaffold scaffold split.
Figure 6: Fine-tuning performance of the foundation model pre-trained with different sizes of data.
We also conduct experiments to investigate how the sizes of the pre-training and downstream datasets affect performance on downstream tasks. As a case study, we use HOMO-LUMO gap prediction on the Molecule3D dataset. First, we evaluate the effect of downstream dataset size by fine-tuning the same foundation model on varying amounts of downstream data. The results, shown in Fig. 5, demonstrate that increasing the amount of downstream data consistently improves performance for both the random and scaffold splits. More importantly, the performance gain from fine-tuning the MLIP foundation model over training from scratch is especially pronounced when the downstream data is limited. This suggests that the foundation models is particularly beneficial in low-data regimes for downstream tasks.
Table 7: Comparison between fine-tuning the foundation model and training from scratch using SchNet as the backbone on the Molecule3D dataset.
@@
Second, we assess how pre-training dataset size impacts downstream performance by fine-tuning the MLIP foundation model pre-trained with different amounts of data while keeping the downstream data fixed. As shown in Fig. 6, downstream accuracy improves steadily as the size of the pre-training dataset increases. These results highlight the dual importance of both pre-training scale and downstream data availability, and underscore the value of foundation models in resource-constrained settings. Additionally, we pre-train a different model architecture to demonstrate that the benefits of pre-training generalize across architectures. Specifically, we pre-train SchNet as the foundation model and observe that it also outperforms training from scratch, as shown in Table 7, which further validates the effectiveness of the MLIP pre-training.

### v2

@@ 3 Related Work
Molecular Representation Learning. Learning informative molecular representations is critical for accurately predicting molecular properties. Early efforts focused on pre-training models on SMILES strings using language modeling techniques such as BERT (Devlin et al., 2019), exemplified by SMILES-BERT (Wang et al., 2019). Subsequently, attention shifted toward pre-training on 2D molecular graphs by designing self-supervised learning tasks (Hu et al., 2019; Rong et al., 2020). Other works further explored learning shared representations between 2D and 3D molecular graphs through contrastive learning (Liu et al., 2021; Stärk et al., 2022). More recently, denoising pre-training (Zaidi et al., 2022; Feng et al., 2023; Ni et al., 2023; Liao et al., 2024) has emerged as a highly effective approach for molecular representation learning. In this paradigm, 3D molecular structures are perturbed with specific noise, and the model is trained to predict the applied noise. Denoising pre-training has been shown to outperform previous pre-training methods, as it is mathematically equivalent to learning an underlying interatomic potential, leading to richer and more physically grounded representations. Another line of work focuses on 3D GNNs that learn informative representations of molecular geometries to predict molecular properties. Representative methods include SchNet (Schütt et al., 2017), SphereNet (Liu et al., 2022), ComENet (Wang et al., 2022), DimeNet++ (Gasteiger et al., 2020), TorchMD-Net (Thölke & De Fabritiis, 2022), and PaiNN (Schütt et al., 2021), etc. A key limitation of these methods is that they require access to ground-truth 3D geometries in order to achieve accurate predictions.
Molecular Foundation Models. Inspired by the remarkable success of pre-trained models in computer vision (Bommasani et al., 2021; Liu et al., 2024; Zhai et al., 2022; Dehghani et al., 2023) and natural language processing (Brown et al., 2020; Touvron et al., 2023; Achiam et al., 2023; Bai et al., 2023), molecular pre-trained models have also begun to attract significant attention. However, due to the lack of large-scale 3D molecule data with energy and force labels, existing molecule pre-trained models primarily focus on 1D SMILES strings or 2D molecular graph representations. For example, ChemFM (Cai et al., 2024) adopts self-supervised causal language modeling on SMILES, while MolE (Méndez-Lucio et al., 2022) adapts DeBERTa (He et al., 2020) for molecular graphs. MolFM (Luo et al., 2023) jointly learns molecular representations from graphs, biomedical texts, and knowledge graphs. Nevertheless, none of these models address the fundamental task of learning 3D molecular energies and forces, which limits their applicability to geometry-dependent tasks such as 3D molecular property prediction, geometry optimization, etc.
@@ 4 Experiments
In this section, we demonstrate the effectiveness of the MLIP pre-trained model across several downstream tasks. In Section˜4.1, we benchmark several backbone model candidates on PubChemQCR-S. In Section˜4.2, we present geometry optimization using the pre-trained model. In Section˜4.3, we show that pre-trained model-relaxed geometries can improve molecular property prediction when stable 3D structures are unavailable in the test set. In Section˜4.4, we fine-tune the pre-trained model for 3D molecular property prediction.
@@
Model Energy MAE (meV/atom) Force RMSE (meV/Å) Time (min/epoch)
SchNet $5.30$ $56.55$ $25$
PaiNN $5.13$ $46.34$ $26$
NequIP $7.37$ $54.78$ $130$
SevenNet $8.77$ $47.63$ $150$
Allegro $10.86$ $60.71$ $85$
FAENet $7.28$ $60.24$ $16$
MACE $7.54$ $51.46$ $120$
PACE $6.24$ $50.54$ $140$
Equiformer $\mathbf{4.69}$ $\mathbf{34.67}$ $65$
@@ 4.1 Backbone Models
To select the backbone architecture for the MLIP pre-trained model, we benchmark several representative MLIP models on our curated dataset. The methods include SchNet (Schütt et al., 2018), FAENet (Duval et al., 2023), NequIP (Batzner et al., 2022), Equiformer (Liao & Smidt, 2022), SevenNet (Park et al., 2024), Allegro (Musaelian et al., 2023), PaiNN (Schütt et al., 2021), PACE (Xu et al., 2024), and MACE (Batatia et al., 2022).
For training efficiency, we curated a smaller subset for model benchmarking, named PubChemQCR-S, which contains 40,979 trajectories and 1,504,431 molecular snapshots from the DFT stage calculations. The benchmark results are summarized in Table˜2. In selecting a pre-trained model backbone, we consider both predictive performance and computational efficiency. Among the candidates, PaiNN demonstrates relatively strong energy prediction accuracy and the second-best force prediction performance on the benchmark while maintaining computational efficiency, making it a strong choice for large-scale pre-training. Detailed introduction and training of those benchmarked methods can be found in Appendices˜B and C.
@@ 4.2 Geometry Optimization
Table 3: Geometry optimization performance of the MLIP pre-trained model pre-trained on the curated PubChemQCR dataset.
@@
Metrics. Since the pre-trained model is trained using data from the DFT relaxation stage, we evaluate its geometry optimization performance by relaxing molecules starting from the first snapshot of the DFT trajectory. The evaluation metrics we adopt are partially based on those proposed in (Tsypin et al., 2023), and include: (1) Average Energy Minimization Percentage, $\overline{\mathrm{pct}}_{T}$, quantifies how much energy is minimized by the MLIP-based optimization relative to the DFT-based optimization; (2) Chemical Accuracy Success Rate, $\mathrm{pct}_{\text{success}}$, measures the percentage of relaxed molecules whose residual energy is within chemical accuracy (commonly defined as 1 kcal/mol); (3) Divergence Rate, $\mathrm{pct}_{\text{div}}$, represents the percentage of relaxed molecules for which either the single-point DFT energy calculation failed or the relaxed DFT energy is higher than the initial energy; (4) Force Convergence Rate, $\mathrm{pct}_{\text{FwT}}$, measures the percentage of relaxed molecules whose maximum force is below a threshold of 0.05 eV/Å. More details can be found in Appendix˜D.
Results. The results are presented in Table˜3. While the value of $\overline{\mathrm{pct}}_{T}$ indicates that the MLIP-relaxed geometries reduce a certain amount of energy compared to the initial structures, a notable portion of conformers remains outside the optimal region. Consequently, the values of $\mathrm{pct}_{\text{success}}$ and $\mathrm{pct}_{\text{FwT}}$ are relatively low. This highlights the inherent challenge of optimizing molecular geometries that are already near the energy minimum—a regime where achieving further relaxation requires extremely accurate modeling of the potential energy surface.
@@
4.3 Molecular Property Prediction with MLIP Pre-Trained Model Relaxed Geometries
Task. As discussed in Section˜1, obtaining accurate 3D geometries is crucial for achieving strong performance in quantum property prediction. However, acquiring ground-state geometries typically requires time-consuming DFT-based relaxation. To address this limitation, it is highly desirable to leverage machine learning interatomic potentials (MLIPs) to generate approximate stable geometries that can support downstream property prediction. In this task, we focus on a practical setting where ground-truth 3D geometries are available during training, but only non-stable geometries are available at test time—consistent with the setup used in Uni-Mol+.
Dataset. We use the Molecule3D dataset and predict the HOMO-LUMO gap, a key quantum property of molecular electronic structure. We use a subset of the dataset containing 600,000 molecules for random and scaffold splits. More details about the dataset can be found in Appendix˜E.
Results. We compare geometry fine-tuning using input structures with varying levels of geometric accuracy. Specifically, we evaluate downstream performance using: (1) RDKit-generated geometries, (2) geometries after the PM3 and HF optimization, i.e. geometries before the DFT relaxation, and (3) MLIP pre-trained model-relaxed geometries. We include (2) because the pre-trained model is trained on DFT substage trajectories and the relaxation begins from the first snapshot of the DFT substage. For a fair comparison with Uni-Mol+, we also evaluate Uni-Mol+ using both RDKit-generated geometries and geometries after the PM3 and HF optimization. The results of geometry fine-tuning on the random and scaffold splits are shown in Table˜4. These results demonstrate that using pre-trained model-relaxed geometries consistently improves downstream property prediction. Furthermore, compared to Uni-Mol+, our approach—combining MLIP-based relaxation with a downstream 3D GNN—achieves superior performance, suggesting that this modular pipeline is more effective than architectures specifically designed to bridge the gap between non-stable and ground truth molecular structures.
@@ 4.3 Molecular Property Prediction with MLIP Pre-Trained Model Relaxed Geometries
4.4 Fine-Tuning MLIP Pre-Trained Model for Molecular Property Prediction
Table 5: Results of HOMO-LUMO gap prediction on the $\nabla^{2}$DFT dataset. Best results are shown in bold, and second-best results are underlined.
@@ 4.4 Fine-Tuning MLIP Pre-Trained Model for Molecular Property Prediction
Task. To demonstrate the effectiveness of the MLIP pre-trained model for 3D molecular property prediction, we fine-tune the pre-trained pre-trained model on downstream tasks where the goal is to predict molecular properties given 3D molecular structures. In this setting, ground-truth geometries are used as input. We evaluate the performance on two benchmark datasets, as described below.
Datasets. We use Molecule3D subset created in Section˜4.3 and $\nabla^{2}$DFT (Khrabrov et al., 2024) datasets. The details of the datasets can be found in Appendix˜E. In this task, we predict the HOMO-LUMO gap, and ground truth 3D geometries are provided. Additional results on the full Molecule3D dataset are provided in Section˜G.1.
Baselines. We select a set of representative models as baselines, including GIN-virtual (Hu et al., 2021), SchNet (Schütt et al., 2017), SphereNet (Liu et al., 2022), ComENet (Wang et al., 2022), DimeNet++ (Gasteiger et al., 2020), TorchMD-Net (Thölke & De Fabritiis, 2022), Uni-Mol+ (Lu et al., 2023), and PaiNN (Schütt et al., 2021). For the Molecule3D dataset, we include 2D models, 3D models, and hybrid approaches like Uni-Mol+. For the $\nabla^{2}$DFT dataset, we focus exclusively on 3D models.
@@
Results. The results for the Molecule3D random split and scaffold split are reported in Table˜6. The fine-tuned pre-trained model achieves the best performance in both settings. The scaffold split is a more challenging evaluation setup, as it requires models to generalize to out-of-distribution molecular scaffolds. As expected, all methods perform worse on the scaffold split compared to the random split, and the gap between validation and test accuracy is also larger in the scaffold setting. GIN-virtual performs the worst among all methods, as it only uses the 2D molecular graph as input, while the HOMO-LUMO gap is highly sensitive to 3D molecular geometry. Uni-Mol+ outperforms GIN-virtual by incorporating ground-truth 3D geometries during training and learning to predict stable 3D structures from RDKit-initialized conformers. However, it still lags behind 3D GNN models, which use accurate 3D geometries during both training and inference. The results on the $\nabla^{2}$DFT dataset are presented in Table˜5. Again, the fine-tuned pre-trained model achieves the best performance among the 3D baselines, further demonstrating its effectiveness in downstream molecular quantum property prediction tasks.
Table 6: Results of HOMO-LUMO gap prediction on two splits of the Molecule3D dataset. Force2Prop w/ PaiNN denotes the MLIP model pre-trained on our PubChemQCR dataset. Best results are shown in bold, and second-best results are underlined.
@@
Figure 5: Fine-tuning performance of the pre-trained model pre-trained with different sizes of data.
Analysis. Inspired by Zaidi et al. (2022), we conduct an experiment to evaluate the usefulness of pre-trained features. Specifically, we freeze the backbone and fine-tune only the prediction head, and compare it to a baseline where the backbone is randomly initialized and only the prediction head is trained. The training curves are shown in Fig.˜4. As expected, fine-tuning only the prediction head performs worse than fine-tuning the entire model, but still significantly outperforms training the head on top of random features. This indicates that pre-trained features provide a meaningful representation for the downstream task. However, to fully adapt the model and achieve optimal performance, it is necessary to fine-tune the entire network.
Figure 6: Results of fine-tuning the MLIP pre-trained model on Molecule3D with different sizes of downstream data for LABEL:sub@fig:different_downstream_data_size_random random and LABEL:sub@fig:different_downstream_data_size_scaffold scaffold split.
We also conduct experiments to investigate how the sizes of the pre-training and downstream datasets affect performance on downstream tasks. As a case study, we use HOMO-LUMO gap prediction on the Molecule3D dataset. First, we evaluate the effect of downstream dataset size by fine-tuning the same pre-trained model on varying amounts of downstream data. The results, shown in Fig.˜6, demonstrate that increasing the amount of downstream data consistently improves performance for both the random and scaffold splits. More importantly, the performance gain from fine-tuning the MLIP pre-trained model over training from scratch is especially pronounced when the downstream data is limited. This suggests that the pre-trained models is particularly beneficial in low-data regimes for downstream tasks.
Table 7: Comparison between fine-tuning the pre-trained model and training from scratch using SchNet as the backbone on the Molecule3D dataset.
@@
Second, we assess how pre-training dataset size impacts downstream performance by fine-tuning the MLIP pre-trained model pre-trained with different amounts of data while keeping the downstream data fixed. As shown in Fig.˜5, downstream accuracy improves steadily as the size of the pre-training dataset increases. These results highlight the dual importance of both pre-training scale and downstream data availability, and underscore the value of pre-trained models in resource-constrained settings. Additionally, we pre-train a different model architecture to demonstrate that the benefits of pre-training generalize across architectures. Specifically, we pre-train SchNet as the pre-trained model and observe that it also outperforms training from scratch, as shown in Table˜7, which further validates the effectiveness of the MLIP pre-training.